# CTNeRF: Cross-Time Transformer for Dynamic Neural Radiance Field from Monocular Video

### v1

@@ Abstract
The goal of our work is to generate high-quality novel views from monocular videos of complex and dynamic scenes. Prior methods, such as DynamicNeRF, have shown impressive performance by leveraging time-varying dynamic radiation fields. However, these methods have limitations when it comes to accurately modeling the motion of complex objects, which can lead to inaccurate and blurry renderings of details. To address this limitation, we propose a novel approach that builds upon a recent generalization NeRF, which aggregates nearby views onto new viewpoints. However, such methods are typically only effective for static scenes. To overcome this challenge, we introduce a module that operates in both the time and frequency domains to aggregate the features of object motion. This allows us to learn the relationship between frames and generate higher-quality images. Our experiments demonstrate significant improvements over state-of-the-art methods on dynamic scene datasets. Specifically, our approach outperforms existing methods in terms of both the accuracy and visual quality of the synthesized views.
Index Terms:
@@
publicationid: pubid: Copyright ©20xx IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
I Introduction
@@
Realistically rendering and presenting dynamic real-world scenes is a highly challenging research topic, with diverse applications in fields such as film production and virtual reality [1, 2, 3]. However, accurately modeling these scenes using traditional mesh-based methods can be difficult due to the complex movements of multiple objects and changes in factors like mirroring and transparency that occur during these movements. While multi-view-based methods have shown better results, they come with their own limitations. These methods require a large number of cameras, resulting in high costs and technical challenges like synchronization and data processing [2, 4, 5, 6, 7]. Additionally, they are not easily applicable in daily life scenarios. Although reconstruction from monocular videos is a promising approach for scene reconstruction, novel view synthesis for monocular videos of dynamic scenes is a more challenging.
@@
Recent advancements in deep learning have made significant breakthroughs in novel view synthesis, with Neural Radiative Fields (NeRF) [8, 9, 10, 11, 12, 13, 14] being one of the most notable contributions to this area. NeRF employs the position and viewing direction of a given image as a query, and employs volume rendering to generate the color of each pixel. However, these methods are primarily designed for static scenes and do not perform optimally when dealing with dynamic objects or scenes. To address this limitation, recent research has explored the application of this approach to monocular dynamic video [15, 16, 17, 18, 19, 20]. For example, some studies have focused on learning a deformable warp field [21, 22] or a neural scene flow between adjacent frames [17, 20, 16, 23]. These efforts aim to extend the utility of NeRF and enable more robust and accurate synthesis of dynamic scenes. Despite the success achieved by NeRF-based methods for dynamic scenes, they still have some limitations. For instance, deformable warp field methods such as Nerfies [22] can handle long sequences but may not perform well for dynamic scenes with complex object motion. On the other hand, neural scene flow or neural trajectory methods like NSFF [17] can handle large movements in dynamic scenes, but their effectiveness is highly dependent on the accuracy of the predicted scene flow or trajectory.
@@
We propose a novel approach that can be applied to dynamic scenes, enabling the handling of more complex motions and improving the rendering results. Our method draws inspiration from recent research on rendering static scenes [14, 24, 25, 26, 27], where local image features are synthesized by aggregating them along epipolar lines from nearby views to enhance the rendering process. However, the apparent limitations assumed by these methods are violated by scenes in motion, making them unsuitable for direct application to dynamic scenes. To overcome this challenge, we have designed a module that aggregates changes in light due to motion in the light space, along with the obtained multi-view features. This enables us to accurately consider both temporal and spatial changes in geometry and appearance, resulting in better rendering of dynamic scenes. More specifically, we first input the extracted feature vector into a cross-time transformer. Next, we input the feature aggregated with time information into a ray transformer to find the relationship between the sampling points on the ray and obtain the aggregated feature. In addition, to strengthen the spatial-temporal relationship of feature vectors, we use a 2D fast Fourier transform frequency-domain feature aggregation module to obtain the aggregated features. Finally, we feed the fused feature vectors of these two features together with the queried rays into residual-based MLPs to output color and density. Experimental results demonstrate that our method can synthesize new views with high quality. Furthermore, compared to previous methods, our approach can render higher-quality ground-truth details of ground truth in dynamic regions. In summary, the contributions of our work are as follows:
@@
II Related Work
II-A Novel view synthesis
Image-based novel view synthesis methods typically leverage input images to integrate pixel information, often benefiting from multi-view approaches for enhanced performance. Techniques like light field [28, 29, 30] and Lomography rendering utilize filtering and interpolating ray samples to generate novel views without relying on an explicit geometric model. However, these methods demand capturing and storing numerous views for optimal rendering [31, 32, 33, 34, 35]. Several approaches employ explicit proxy geometry to achieve high-quality rendering with only a few input images [36, 37, 38, 39, 40, 32, 41, 42, 43, 44]. However, challenges such as untextured regions, highlights, reflections, and repetitive patterns may lead to difficulties in accurately estimating scene geometry. To tackle these issues, prior works have explored techniques such as local warps, soft 3D reconstruction, and learning-based approaches. To overcome the aforementioned issues, Zhang et al. [45] proposed an end-to-end solution for single-image-based view synthesis, aiming for maximum deformability without the need for deep priors. They initially transform the input image into a latent feature space and then reconstruct new views in the image domain. Although this approach can optimize errors and artifacts in the synthesized views, challenges persist in handling complex texture regions.
II-B Dynamic region view synthesis
Dynamic region novel view synthesis is a challenging task because it involves generating a new view of a scene from arbitrary viewpoints that may not be presented in the original input data [48, 49]. This requires the model to understand the scene’s geometry, lighting, and texture and to extrapolate this information to generate realistic new views. However, most previous work for dynamic region view synthesis requires multi-view video as input [50, 51, 7, 52, 53, 54, 55, 56, 5, 57], which may not be applicable or practical in real-world scenarios. This limitation can make it difficult to obtain the necessary data to train the model effectively and generate high-quality synthesized views. To address this limitation, some methods have been proposed that use a single camera for the synthesis of the dynamic region. Yoon et al. [58] propose a method that leverages warping from a depth map to synthesize novel views. They use multi-view stereo for static regions and monocular depth estimation for dynamic regions and then blended for warping to the final result. However, this approach is highly dependent on the accuracy of the depth estimation and may lead to errors and artifacts in the synthesized view when the scene contains occlusions or reflections.
With NeRF [8, 9, 10, 11, 12, 13, 14] demonstrating impressive results in view synthesis tasks, several works have attempted to extend NeRF to tackle dynamic new view synthesis challenges [16, 20, 17, 18]. These methods can be classified into two main directions. The first direction involves using deformation fields to represent scenes [18, 21, 22, 59]. While this approach can handle long sequences of videos, its primary challenge is dealing with large motions in the scene. As this method typically warps the scene from the same frame, it can result in a lack of continuity throughout the entire sequence, such as Nerfies [22] and HyperNeRF [21]. The second approach is based on the time-varying 4D radiance fields approach [16, 60, 17, 61, 10]. These methods model dynamic scenes as time-varying continuous functions of appearance, geometry, and 3D scene motion by predicting the scene flow field. Although such methods can capture fast and complex motion in the scene, they usually require more accurate scene flow or trajectory guidance and cannot handle non-rigid deformation well. Our proposed method aggregates feature from nearby views to effectively handle this situation and improve rendering results.
III Proposed Method
In this section, we present the proposed methods with the goal of enabling the trained model to query new viewpoints at any time and angle within a monocular video of a dynamic scene. Our system pipeline (Figure 1) can be divided into three parts: Each frame $I_{i}$ of the video corresponds to a known camera parameter $P_{i}$. Similar to other time-varying NeRF-based techniques, we first optimize the model to reconstruct the input frame, before being utilized for rendering novel views. Instead of directly encoding 3D color and density in the weights of the MLPs like recent dynamic NeRF methods [16, 17], we borrow the idea of a recent generalized NeRF to aggregate features from views near the target view to enhance rendering. Below we describe our approach to multi-feature aggregation, a ray-based cross-time aggregation module, and a boost module via frequency-domain effects.
III-A Multi-view aggregation
@@
III-A1 Static Region Feature Aggregation
@@
$x_{t\pm i}=P_{\pm i}x_{t}$ (1)
where $x_{t}$ is a point in space on the camera ray on the target view, the $x_{t\pm i}$ is the adjacent view, and the $P_{\pm i}=[R_{\pm i},T_{\pm i}]$ is the camera parameters, note that we set $i=1$. And the queried feature can be expressed as:
$F_{t\pm i}=E(proj\left\langle x_{t\pm i}\right\rangle)$ (2)
$proj\left\langle\cdot\right\rangle$ represents the coordinates of the point projection image in space, and then uses a feature extractor $E(\cdot)$ to extract the features of the image, and finally obtains the query feature vector $F_{t\pm i}$. The RGB $c$ and density $\sigma$ of the static region can use an MLP to query, which can be expressed as:
@@
the inputs include extracted feature vectors $F_{t\pm i}$, target view direction $d_{t}$, and space coordinates $x,y,z$.
@@
III-A2 Dynamic Region Feature Aggregation
@@
$x_{t+1}=x_{t}+s_{fw}$ (4)
$x_{t-1}=x_{t}+s_{bw}$ (5)
where, $s_{fw}$ and $s_{bw}$ are the predicted scene flow. And then, we can obtain the corresponding feature vectors using Equation 2. Note that we use a four-layer MLP to predict this scene flow, while also outputting a blend value to weigh the color and density of static and dynamic regions. Thus, it can be expressed as:
@@
$s_{fw}$ and $s_{bw}$ are the predicted scene flow, $b$ is the predicted blending value and the $i={\{0,1\}}$.
III-A3 Combining static and dynamic models
@@
III-B Ray-Based Cross-Time Aggregation Module
III-B1 Cross-Time Aggregation
By means of the aforementioned method, the correlation between the frontal and posterior frames can be integrated using the camera rays from the current viewing angle. However, it should be noted that there exists a temporal relationship among the input feature vectors of adjacent frames, and a direct aggregation would overlook this temporal variation. Therefore, we design a cross-time converter to enable the feature vector of the target frame to attentively focus on the variations in the feature vector of its adjacent frame (as illustrated in the left figure, see Figure 2). The cross-time transformer leverages a classic cross-attention mechanism [63, 64] that allows for a variable number of inputs, given the frame-to-frame changes involved:
@@ III-B1 Cross-Time Aggregation
TABLE I: Novel view synthesis quantitative results on Nvidia Dynamic Scene Dataset [58]. We report average PSNR and LPIPS [65] results by comparison with existing methods. The best performance is bold and the next best is underline.
@@
TABLE II: Assessing novel view synthesis outcomes, we measure performance using the mPSNR and mSSIM metrics, benchmarked against established methods. The evaluation is conducted on the iPhone dataset [60].

### v2

@@ Abstract
The goal of our work is to generate high-quality novel views from monocular videos of complex and dynamic scenes. Prior methods, such as DynamicNeRF, have shown impressive performance by leveraging time-varying dynamic radiation fields. However, these methods have limitations when it comes to accurately modeling the motion of complex objects, which can lead to inaccurate and blurry renderings of details. To address this limitation, we propose a novel approach that builds upon a recent generalization NeRF, which aggregates nearby views onto new viewpoints. However, such methods are typically only effective for static scenes. To overcome this challenge, we introduce a module that operates in both the time and frequency domains to aggregate the features of object motion. This allows us to learn the relationship between frames and generate higher-quality images. Our experiments demonstrate significant improvements over state-of-the-art methods on dynamic scene datasets. Specifically, our approach outperforms existing methods in terms of both the accuracy and visual quality of the synthesized views. Our code is available on https://github.com/xingy038/CTNeRF.
keywords:
@@
journal: Pattern Recognition
1 Introduction
@@
Realistically rendering and presenting dynamic real-world scenes is a highly challenging research topic, with diverse applications in fields such as film production and virtual reality [1, 2, 3]. However, accurately modeling these scenes using traditional mesh-based methods can be difficult due to the complex movements of multiple objects and changes in factors like mirroring and transparency that occur during these movements. While multi-view-based methods have shown better results, they come with their own limitations. These methods require a large number of cameras, resulting in high costs and technical challenges like synchronization and data processing [4, 5, 6, 7]. Additionally, they are not easily applicable in daily life scenarios. Although reconstruction from monocular videos is a promising approach for scene reconstruction, novel view synthesis for monocular videos of dynamic scenes is more challenging.
@@
Recent advancements in deep learning have made significant breakthroughs in novel view synthesis, with Neural Radiative Fields (NeRF) [8, 9] being one of the most notable contributions to this area. NeRF employs the position and viewing direction of a given image as a query, and employs volume rendering to generate the color of each pixel. However, these methods are primarily designed for static scenes and do not perform optimally when dealing with dynamic objects or scenes. To address this limitation, recent research has explored the application of this approach to monocular dynamic video [10, 11]. For example, some studies have focused on learning a deformable warp field [12] or a neural scene flow between adjacent frames [13, 14, 11, 15]. These efforts aim to extend the utility of NeRF and enable more robust and accurate synthesis of dynamic scenes. Despite the success achieved by NeRF-based methods for dynamic scenes, they still have some limitations. For instance, deformable warp field methods such as Nerfies [16] can handle long sequences but may not perform well for dynamic scenes with complex object motion. On the other hand, neural scene flow or neural trajectory methods like NSFF [13] can handle large movements in dynamic scenes, but their effectiveness is highly dependent on the accuracy of the predicted scene flow or trajectory.
@@
We propose a novel approach that can be applied to dynamic scenes, enabling the handling of more complex motions and improving the rendering results. Our method draws inspiration from recent research on rendering static scenes [17, 18, 19, 20], where local image features are synthesized by aggregating them along epipolar lines from nearby views to enhance the rendering process. However, the apparent limitations assumed by these methods are violated by scenes in motion, making them unsuitable for direct application to dynamic scenes. To overcome this challenge, we have designed a module that aggregates changes in ray due to motion in the ray space, along with the obtained multi-view features [21]. This enables us to accurately consider both temporal and spatial changes in geometry and appearance, resulting in better rendering of dynamic scenes. More specifically, we first input the extracted feature vector into a cross-time transformer. Next, we input the feature aggregated with time information into a ray transformer to find the relationship between the sampling points on the ray and obtain the aggregated feature. In addition, to strengthen the spatial-temporal relationship of feature vectors, we use a 2D fast Fourier transform frequency-domain feature aggregation module to obtain the aggregated features. Finally, we feed the fused feature vectors of these two features together with the queried rays into residual-based MLPs to output color and density. Experimental results demonstrate that our method can synthesize new views with high quality. Furthermore, compared to previous methods, our approach can render higher-quality ground-truth details of ground truth in dynamic regions. In summary, the contributions of our work are as follows:
@@
2 Related Work
2.1 Novel view synthesis
2.2 Dynamic region view synthesis
With NeRF [22, 8, 9] demonstrating impressive results in view synthesis tasks, several works have attempted to extend NeRF to tackle dynamic new view synthesis challenges [11, 14, 13, 29]. These methods can be classified into two main directions. The first direction involves using deformation fields to represent scenes [29, 12, 16, 30]. While this approach can handle long sequences of videos, its primary challenge is dealing with large motions in the scene. As this method typically warps the scene from the same frame, it can result in a lack of continuity throughout the entire sequence, such as Nerfies [16] and HyperNeRF [12]. The second approach is based on the time-varying 4D radiance fields approach [31, 13, 32]. These methods model dynamic scenes as time-varying continuous functions of appearance, geometry, and 3D scene motion by predicting the scene flow field. Although such methods can capture fast and complex motion in the scene, they usually require more accurate scene flow or trajectory guidance and cannot handle non-rigid deformation well. Our proposed method aggregates feature from nearby views to effectively handle this situation and improve rendering results.
@@ 2.2 Dynamic region view synthesis
3 Proposed Method
In this section, we present the proposed methods with the goal of enabling the trained model to query new viewpoints at any time and angle within a monocular video of a dynamic scene. Our system pipeline (Figure 1) can be divided into two parts: one part focuses on the static background, while the other part handles the dynamic foreground, and finally blends the two through blending to obtain the reconstructed video. Furthermore, similar to other time-varying NeRF-based techniques, we first optimize the model to reconstruct the input frame, before being utilized for rendering novel views. Instead of directly encoding 3D color and density in the weights of the MLPs like recent dynamic NeRF methods [11, 13], we borrow the idea of a recent generalized NeRF to aggregate features from views near the target view to enhance rendering. Below we describe our approach to multi-feature aggregation, a ray-based cross-time aggregation module, and a boost module via frequency-domain effects.
3.1 Multi-view aggregation
@@
3.1.1 Static Region Feature Aggregation
@@
$x_{t\pm i}=P_{\pm i}x_{t}\in\mathbb{R}^{3}$ (1)
where $x_{t}$ is a point in space on the camera ray on the target view, the $x_{t\pm i}\in\mathbb{R}^{3}$ is the adjacent view, and the $P_{\pm i}=[R_{\pm i},T_{\pm i}]\in\mathbb{R}^{4\times 4}$ is the camera parameters, note that we set $i=1$. And the queried feature can be expressed as:
$F_{t\pm i}=E(proj\left\langle x_{t\pm i}\right\rangle)\in\mathbb{R}^{d}$ (2)
$proj\left\langle\cdot\right\rangle$ represents the coordinates of the point projection image in space, and then uses a feature extractor $E(\cdot)$ to extract the features of the image, and finally obtains the query feature vector $F_{t\pm i}\in\mathbb{R}^{d}$. The RGB $c$ and density $\sigma$ of the static region can use an MLP to query, which can be expressed as:
@@
the inputs include extracted feature vectors $F_{t\pm i}\in\mathbb{R}^{d}$, target view direction $d_{t}\in\mathbb{R}^{3}$, and space coordinates $x,y,z$.
@@
3.1.2 Dynamic Region Feature Aggregation
@@
$x_{t+1}=x_{t}+s_{fw}\in\mathbb{R}^{3}$ (4)
$x_{t-1}=x_{t}+s_{bw}\in\mathbb{R}^{3}$ (5)
where, $s_{fw}\in\mathbb{R}^{3}$ and $s_{bw}\in\mathbb{R}^{3}$ are the predicted scene flow. And then, we can obtain the corresponding feature vectors using Equation 2. Note that we use a four-layer MLP to predict this scene flow, while also outputting a blend value to weigh the color and density of static and dynamic regions. Thus, it can be expressed as:
@@
$s_{fw}\in\mathbb{R}^{3}$ and $s_{bw}\in\mathbb{R}^{3}$ are the predicted scene flow, $b$ is the predicted blending value and the $i={\{0,1\}}$.
3.1.3 Combining static and dynamic models
@@
3.2 Ray-Based Cross-Time Aggregation Module
3.2.1 Cross-Time Aggregation
By means of the aforementioned method, the correlation between the frontal and posterior frames can be integrated using the camera rays from the current viewing angle. However, it should be noted that there exists a temporal relationship among the input feature vectors of adjacent frames, and a direct aggregation would overlook this temporal variation. Therefore, we design a cross-time converter to enable the feature vector of the target frame to attentively focus on the variations in the feature vector of its adjacent frame (as illustrated in the left figure, see Figure 3). The cross-time transformer leverages a classic cross-attention mechanism [33, 34, 35, 36] that allows for a variable number of inputs, given the frame-to-frame changes involved:
@@ 3.2.1 Cross-Time Aggregation
3.2.2 Ray Aggregation
=Ray $\mathcal{F}=\text{Ray\ Transformer}(\hat{F}_{t\pm i})$ (15)
Specifically, we define:
$\mathcal{F}=Softmax(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}})\cdot V$ (17)
3.3 Frequency Domain Aggregation Module
We introduce a novel spatio-temporal feature learning module termed the Global Spatio-Temporal Filter (GSTF), inspired by recent advancements in frequency-domain-based methodologies [38] aimed at enhancing the rendering quality of new views [39]. The primary objective of GSTF is to elevate the representation of feature vectors by capturing both spatial and temporal relationships through specialized frequency filters. In our approach, GSTF is meticulously crafted to discern and learn distinct frequency filters at each spatial location. This enables the modeling of temporal variations within feature vectors across different spatial positions. The core mechanism involves the transformation of both temporal and spatial features at each location into frequency feature spectra. This transformation is achieved through a two-dimensional Fast Fourier Transform (FFT) [40]. The frequency filter, learned through GSTF, acts as a modulator on this transformed spectrum. Subsequently, we revert this modulated spectrum back to the time domain using an inverse FFT. This comprehensive process allows GSTF to effectively encode the intricate interplay between time and space in the feature vectors, contributing to the improvement of new view rendering. To gain a better understanding of our GSTF design, let’s first review the convolution theorem in the field of digital signal processing [41]. Given a sequence of feature signals with $T$ points ($f[t],0\leq t\leq T-1$), we can calculate its discrete spectrum $S[k]$ using Discrete Fourier Transform (DFT) via:
$S[k]=\sum_{t=0}^{T-1}f[t]e^{-j(2\pi/T)kt},0\leq k\leq T-1$ (18)
$f[t]=\frac{1}{T}\sum_{k=0}^{T-1}S[k]e^{j(2\pi/T)kt},0\leq t\leq T-1$ (19)
Specifically, we first convert the features into frequency domain signals. Next, these frequency domain signals are filtered, and then the filtered signals are reconstructed into time domain features. Our GSTF can be easily used in modern deep-learning frameworks such as PyTorch [42], The pseudocode of PyTorch is shown in Algorithm 1. Finally, these time-domain features are merged with the RBCT-processed features to achieve effective aggregation of time-domain and frequency-domain features. It can be expressed as:
$F=f[t]+\mathcal{F}$ (20)
Where $f[t]$ is the frequency-domain features, $\mathcal{F}$ is the time-domain features, and $F$ is the aggregated features. Through ablation experiments, we demonstrate that our global filtering mechanism is an effective spatial information mixing method. The GSTF module improves the texture quality of new views and alleviates blurring and artifacts often observed in new view synthesis.
Algorithm 1 Global Spatio-Temporal Filter
1:Initialization: learnable weight w
2:x = torch.fft.rfft2(x, dim=(0,1))
3:x = x * w
4:x = torch.fft.irfft2(x, dim=(0,1))
3.4 Regularization
$\mathcal{L}=\mathcal{L}_{data}+\mathcal{L}_{small}+\mathcal{L}_{pho}$ (21)
3.5 Architecture
Our approach utilizes two distinct types of models: static and dynamic. The architecture of static and dynamic model is depicted in Figure 4.
Figure 4: Network architectures of our static and dynamic representations.
4 Experiment
4.1 Implementation details
4.2 Model Parameters and Inference Time
Method Parameters Inference speed Training Time
DynamicNeRF 4.59M 8 s/Frame about 21 hours
Our 12.21M 10 s/Frame about 24 hours
Table 1: Parameters, inference speed, and training time.
We present our model’s Parameters, inference speed, and training time. The amount of our speed parameters has increased compared to our baseline, but in more complex scenes, our baseline cannot correctly obtain dynamic scenes (As shown in Figure 6).
4.3 Dataset
Nvidia Dynamic Scene Dataset comprises nine video sequences captured using a static camera rig of 12 cameras. All cameras capture images simultaneously at 12 different time steps $\{t_{0},t_{1},\dots,t_{11}\}$, and we obtain a twelve-frame monocular video $\{I_{0},I_{1},\dots,I_{11}\}$ by sampling the image taken by the $i-th$ camera at time $t_{i}$. It is worth mentioning that we use a different camera for each video frame to simulate camera motion. The video frames consist of a background that remains stationary throughout the video and a dynamic object that changes over time. We adopt COLMAP [46] similar to NeRF [23], to estimate the camera poses, near and far boundaries of the scene, and assume that all cameras share the same intrinsic parameter. We exclude the DynamicFace sequence from our evaluation since COLMAP fails to estimate camera poses for this sequence. Lastly, we resize all video sequences to a resolution of 480 $\times$ 270. The DAVIS Dataset [45] consists of fifty sequences featuring dynamic moving objects, like animals and cars. However, due to limitations in camera movement, COLMAP could only estimate camera poses for six out of the fifty sequences, all of which include ground truth object masks. Finally, for the iPhone Dataset [31], we conducted a quantitative evaluation using seven dynamic scenes, each accompanied by ground truth images of novel views.
4.4 Comparison with State-of-the-Art Methods
4.4.1 Qualitative Results
Table 2: Novel view synthesis quantitative results on Nvidia Dynamic Scene Dataset [44]. We report average PSNR and LPIPS [47] results by comparison with existing methods. The best performance is bold and the next best is underline.
@@ 4.4.1 Qualitative Results
4.4.2 Quantitative Results
Table 2 presents the quantitative results obtained from the Nvidia Dynamic Scene Dataset [44]. We adopted the evaluation methodology from DynamicNeRF [11] to synthesize views using the first camera while varying the time on the Nvidia Dynamic Scene Dataset. To evaluate the rendering quality of each method, we employed two widely recognized error metrics: peak signal-to-noise ratio (PSNR) and perceptual similarity (LPIPS) as defined by [47]. Additionally, due to minor differences observed in our ablation study’s results, we incorporated the structural similarity index (SSIM) for a more thorough assessment.
Table 3: Assessing novel view synthesis outcomes, we measure performance using the mPSNR and mSSIM metrics, benchmarked against established methods. The evaluation is conducted on the iPhone dataset [31].